# ND-SDF: Learning Normal Deflection Fields for High-Fidelity Indoor Reconstruction

### v2

@@ B.2 Ablation of adaptive monocular depth loss
Table 10: Quantitative results of the 6 selected scenes from ScanNet++. Our method achieves state-of-the-art performance.
Appendix C More results
C.1 Replica
We present additional visualization results on Replica in Figure 11.
C.2 ScanNet++
We present detailed quantitative results of the selected six scenes from ScanNet++ in Table 10 and provide more visualization results in Figure 12. Compared to the baselines, our method yields significantly improved reconstructions both visually and quantitatively, notably excelling at capturing thin, fine-grained structures while maintaining smoothness in flat regions.
C.3 Tanks and temples
We present additional qualitative comparisons with previous state-of-the-art baselines, including MonoSDF Yu et al. (2022b) and NeuRIS Yu et al. (2022b), as depicted in Figure 13. It is evident that ND-SDF achieves accurate dense surfaces replete with extensive details.
As illustrated in Figure 14 and Figure 15, we further extracted the reconstructed meshes at a resolution of 2048. At this high resolution, the meshes demonstrate that our method significantly enhances the generation of small and thin structures, even in challenging large-scale indoor scenes. For all experiments, the hash grid settings were kept at their defaults, with resolutions ranging from $2^{5}$ to $2^{11}$ across 16 levels, and a hashmap size of $2^{19}$.
C.4 More information about dataset

### v3

@@ C.3 Tanks and temples
Table 11: Quantitative results of the 6 selected scenes from ScanNet++. ND-SDF significantly outperforms previous state-of-the-art methods in most metrics, particularly in terms of Recall.
C.4 Additional discussions
More information about dataset.
@@ C.4 Additional discussions
Concurrent work.
We notice that the concurrent work NC-SDF Chen et al. (2024b) share analogous idea with us. We clarify that our method is entirely independent of NC-SDF. Although we similarly model an SO(3) residual field to enhance surface reconstruction, our design diverges substantially. NC-SDF employs a two-stage approach, yet struggles with noise in the second stage’s explicit deflection. In contrast, our adaptive prior loss automatically assigns weights based on sample characteristics (Eq. 12), removing the need for staged training. Additionally, NC-SDF requires extra boundary information for optimization, reducing robustness and adding hyperparameters. While NC-SDF was tested on datasets with simple layouts like ScanNet and ICL-NUIM, we evaluated our method on more complex indoor scenes, including ScanNet++ and T&T. Extensive experiments confirm our method’s superior generalization ability in challenging environments, an advantage not observed in NC-SDF.
About unbiasing methods.
It’s crucial to resolve bias issues in SDF-based surface reconstructions. We show the necessity to employ unbiasing transformation through an actual sampling process in Figure 3, supported by ablation results in Figure 5. Inherently, bias refers to the failure of the rendering weight to attain its maximum value at the zero-level set of SDF, thereby leading to incorrect rendering depth. This issue not only damages the surface quality but also severely affects the generation of fine structures. Many works have been proposed to solve this: The unbias assumption of Neus Wang et al. (2021) ensures the local maximum of the weight for a local plane. But the weight is severely reduced when the viewing direction gets closer to the tangent plane. TUVR Zhang et al. (2023) transforms the input SDF value to the actual depth from the tangent plane and extends the extremal property of weight to any surface distribution. But TUVR still generates abnormal bimodal shaped weight when passing close to an object if $1/\beta$ is not large enough. To address this, DebSDF Xiao et al. (2024) proposed using curvature radius to correct the SDF value, tremendously reducing the bias. However, these unbiasing methods, such as TUVR and DebSDF, face serve convergence issues, as the transformation must be perspective-dependent, causing inconsistency. We analyze this issue using TUVR as example in Appendix B.1. To mitigate this, we utilize explicit learned angular deviations to help apply the unbiasing transformation only to low-proportion fine-grained structures in the scene, which has proved effective in our method. Simpler solutions also exist, such as applying unbiasing only after the surface has sufficiently converged. However, another issue arises: the implicit SDF field is difficult to reshape, making it challenging to capture fine structures in later stages. There are still many issues to be addressed.
Appendix D Societal impact
@@
Figure 16: Additional Visualization Comparisons on ScanNet++. We further compare our method with state-of-the-art method DebSDF. As shown above, ND-SDF tremendously outperforms previous methods when facing challenging complex indoor scenes. It shows a superior ability to capture accurate and geometry-consistent detailed structures.
Figure 17: Additional Visualization Comparisons on ScanNet++. We further compare our method with state-of-the-art method DebSDF. As shown above, ND-SDF tremendously outperforms previous methods when facing challenging complex indoor scenes. It shows a superior ability to capture accurate and geometry-consistent detailed structures.